# Tree Posets: Supersaturation, Enumeration, and Randomness

### v2

@@ 4.3 Tools for Induced Posets
$U(\mathcal{S})=\bigcup_{S\in\mathcal{S}}U(S)\text{ and }D(\mathcal{S})=\bigcup_{S\in\mathcal{S}}D(S)$
@@
For every $\delta>0$, and every $\ell,q$ positive integers, there exists $\gamma=\gamma(\delta,\ell,q)$ such that the following holds for every $i,s\in[q]$ such that $i\neq s$. Let $\mathcal{M}$ be a $\ell$-gapped family of $q$-marked chains with markers from ~$\widetilde{\mathcal{B}}_{n}$ and let $F\in\mathcal{L}^{i}(\mathcal{M})$. Suppose $F$ is $(i,s,\gamma)$-bad with respect to $\mathcal{M}$ and that $n$ is sufficiently large.
@@
We will show that $F$ is $(i,\delta)$-lower bound with respect to $\mathcal{M}$ with $W:=\mathcal{W}_{2}\cup D^{*}(F,\mathcal{W}_{1})$ being an $(i,\delta)$-lower witness. Since $\mathcal{M}$ is $\ell$-gapped and for all $D\in\mathcal{W}_{2}$, $D$ is in the $s$th position of some $q$-chain in $\mathcal{M}(F,i)$, we have $|F-D|\geq\ell(s-i)$. Since $n$ is sufficiently large and $F,D\in\widetilde{\mathcal{B}}_{n}$, $|F-D|\leq n/6$. Thus for a uniformly chosen random full chain $\chi\in\mathcal{C}$ we have
@@ 5.1 Embedding Process
Let $P$ be a tree poset of height $k$. Let $q\geq k$ and $\ell$ be positive integers and $\varepsilon>0$ then there exists a constant $\gamma=\gamma(\varepsilon,q,\ell)$ such that the following holds for any $v\in P$ and for any $\mathrm{r}:P\to[q]$ a poset homomorphism. Let ~$\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$ be an $\ell$-gapped family and has Lubell weight $\mu(F)\geq q-1+\varepsilon$ and $\Gamma$ a $(\frac{\gamma}{2},\ell,\mathcal{F})$- bounded family. Then, there is $F\in\mathcal{F}$ such that the number of injective induced homomorphisms $\varphi:P\rightarrow\mathcal{F}^{\prime}$ satisfying $\varphi(P)\not\in\Gamma$ and $\varphi(v)=F$ is at least
@@ 3. $(C_{3})$:
Recall the formal definition of $D^{*}(F,\mathcal{S})$ from Section 4.3 for a member $F\in\mathcal{F}$ and a family $\mathcal{S}$ which is simply the set of all members $S\in\mathcal{S}$ which are downsets of $F$ and are comparable to some set of $\mathcal{S}$. Furthermore, let ^$\widehat{\mathcal{M}}^{|P|-j-1}:=\mathcal{M}^{|P|-j-1}(\varphi^{j}(y),\mathrm{r}(y))$, that is the family of all $q$-marked chains in $\mathcal{M}^{|P|-j-1}$ which have $\varphi^{j}(y)$ in their $\mathrm{r}(y)$th position. Let $\mathcal{P}_{j}=\{\varphi^{j}(z):z\not\geq_{P}y,z\in\{v_{1},v_{2},\dots,v_{j}\}\}$, that is the set of all images of currently embedded elements of $P$ which are either below $y$ or incomparable with $y$, and thus not cannot be comparable with $\varphi^{j+1}(x)$ because in the embedding $\varphi^{j}$, $x$ has the unique neighbor $y$.
@@
For every tree poset $P$ of height $k$, positive integers $q,\ell$ with $q\geq k$, and $\varepsilon>0$, there exists $\gamma=\gamma(\varepsilon,q,\ell)$ such that the following holds for any $\mathrm{r}:P\rightarrow[q]$ a poset homomorphism. Let ~$\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$ be an $\ell$-gapped family such that for all $\mathcal{S}\subseteq\mathcal{F}$ with $|\mathcal{S}|=N$ we have $\mu(\mathcal{F}-\mathcal{S})\geq q-1+\varepsilon$. Then, the number of induced copies of $P$ in $\mathcal{F}$ is at least
@@
Since $F_{1}$ must be a member of the $\mathrm{r}(x_{1})$th layer of $\mathcal{M}_{n,q}$, the number of choices for $F_{1}$ is at most (${n\choose n/2}$. Given that $F_{1},\ldots,F_{j-1}$ have been selected, let $x_{i}$ be the parent of $x_{j}$. If $x_{i}<x_{j}$, then $F_{j}$ must be a set containing $F_{i}$ together with $\mathrm{r}(x_{i})-\mathrm{r}(x_{j})$ additional elements from $[n]$, and the number of such sets is at most $n^{\mathrm{r}(x_{i})-\mathrm{r}(x_{j})}$. Similarly if $x_{j}<x_{i}$ then the number of choices for $F_{j}$ is at most $n^{\mathrm{r}(x_{j})-\mathrm{r}(x_{i})}$. Multiplying the number of choices for each step gives the total result, since each edge of $H(P)$ is counted exactly once by some $n^{|\mathrm{r}(x_{i})-\mathrm{r}(x_{j})|}$ term. ∎
@@
Using Chernoff bounds, it is standard [8, 21] to show that the number of sets $F\in\mathcal{B}_{n}$ with $||F|-n/2|>2\sqrt{n\ln{n}}$ is $o({n\choose n/2})$. Thus, we may assume ~$\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$.
@@ 5.3 Balanced Supersaturation
For every tree poset $P$ with height $k$, real number $\varepsilon>0$, and integer $\ell$; there exists $\delta=\delta(\varepsilon,P,\ell)$ such that the following holds. Let $n$ be sufficiently large and ~$\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$ satisfy $|\mathcal{F}|\geq(k-1+\varepsilon)\binom{n}{n/2}$, and suppose $\mathcal{F}$ is $\ell$-gapped. Then there is a collection $\mathcal{H}$ of induced copies of $P$ from $\mathcal{F}$ satisfying
@@
    $\Delta_{j}(\mathcal{H})\leq(\delta n^{\ell})^{|P|-j}$
    for all $1\leq j\leq|P|$.
@@
    For all $1\leq j\leq|P|$,
    $\Delta_{j}(\mathcal{H})\leq(\delta n^{\ell})^{|P|-\ell},$
then there exists an induced copy $P^{\prime}$ of $P$ not in $\mathcal{H}$ such that =$\mathcal{H}^{\prime}=\mathcal{H}\cup\{P^{\prime}\}$ satisfies
@@ 6.1 Hypergraph Containers
$\tau(\mathcal{F},k):=\begin{cases}\frac{1}{n}&\text{if }|\mathcal{F}|<3k\binom{n}{n/2}\\ \frac{1}{n^{3}}&\text{if }|\mathcal{F}|\geq 3k\binom{n}{n/2}.\end{cases}$
@@
Let $M>0,s>0$ and $10<\delta<$ 1. For any sequence $(a_{1},a_{2},\dots,a_{m})$ of real numbers summing up to $s$ such that $1\leq a_{j}\leq(1-\delta)^{j}M$ for each $j\in[m]$, we have
@@ 6.2 Applications of Container Lemma
Let $k\geq 2$ and $\varepsilon>0$ be arbitrary, and let $K=K(\varepsilon,P)>0$ and $(\mathcal{T},g,C)$ be the constant and fingerprint given by Theorem 6.4 applied to $\mathcal{G}_{P}$ with $\varepsilon,P$. Let $n\in\mathbb{N}$ be sufficiently large, and note that we may assume $pn\geq K\varepsilon^{-1}$ since $pn\rightarrow\infty$. Let $\mathcal{P}(n,p)$ be the uniformly random subset of $\mathcal{B}_{n}$, where each set survives with probability $p$ Suppose $I\subseteq\mathcal{P}(n,p)$ is an induced $P$-free subset (or equivalently an independent set of $\mathcal{G}_{P}$) of size at least $(k-1+3\varepsilon)p\binom{n}{n/2}$. Then, it follows that $\widehat{g(I)}\subseteq\mathcal{P}(n,p)$ and

### v3

@@ 4.3 Tools for Induced Posets
[…]
@@
For every $\delta>0$, and every $\ell,q$ positive integers, there exists $\gamma=\gamma(\delta,\ell,q)$ such that the following holds for every $i,s\in[q]$ such that $i\neq s$. Let $\mathcal{M}$ be a $\ell$-gapped family of $q$-marked chains with markers from $\widetilde{\mathcal{B}}_{n}$ and let $F\in\mathcal{L}^{i}(\mathcal{M})$. Suppose $F$ is $(i,s,\gamma)$-bad with respect to $\mathcal{M}$ and that $n$ is sufficiently large.
@@
We will show that $F$ is $(i,\delta)$-lower bound with respect to $\mathcal{M}$ with $W:=\mathcal{W}_{2}\cup D^{*}(F,\mathcal{W}_{1})$ being an $(i,\delta)$-lower witness. Since $\mathcal{M}$ is $\ell$-gapped and for all $\in\mathcal{W}_{2}, DD$ is in the $s$th position of some $q$-chain in $\mathcal{M}(F,i)$, we have $|F-D|\geq\ell(s-i)$. Since $n$ is sufficiently large and $F,D\in\widetilde{\mathcal{B}}_{n}$, $|F-D|\leq n/6$. Thus for a uniformly chosen random full chain $\chi\in\mathcal{C}$ we have
@@ 5.1 Embedding Process
Let $P$ be a tree poset of height $k$. Let $q\geq k$ and $\ell$ be positive integers and $\varepsilon>0$ then there exists a constant $\gamma=\gamma(\varepsilon,q,\ell)$ such that the following holds for any $v\in P$ and for any $\mathrm{r}:P\to[q]$ a poset homomorphism. Let $\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$ be an $\ell$-gapped family and has Lubell weight $\mu(F)\geq q-1+\varepsilon$ and $\Gamma$ a $(\frac{\gamma}{2},\ell,\mathcal{F})$- bounded family. Then, there is $F\in\mathcal{F}$ such that the number of injective induced homomorphisms $\varphi:P\rightarrow\mathcal{F}^{\prime}$ satisfying $\varphi(P)\not\in\Gamma$ and $\varphi(v)=F$ is at least
@@ 3. $(C_{3})$:
Recall the formal definition of $D^{*}(F,\mathcal{S})$ from Section 4.3 for a member $F\in\mathcal{F}$ and a family $\mathcal{S}$ which is simply the set of all members $S\in\mathcal{S}$ which are downsets of $F$ and are comparable to some set of $\mathcal{S}$. Furthermore, let $\widehat{\mathcal{M}}^{|P|-j-1}:=\mathcal{M}^{|P|-j-1}(\varphi^{j}(y),\mathrm{r}(y))$, that is the family of all $q$-marked chains in $\mathcal{M}^{|P|-j-1}$ which have $\varphi^{j}(y)$ in their $\mathrm{r}(y)$th position. Let $\mathcal{P}_{j}=\{\varphi^{j}(z):z\not\geq_{P}y,z\in\{v_{1},v_{2},\dots,v_{j}\}\}$, that is the set of all images of currently embedded elements of $P$ which are either below $y$ or incomparable with $y$, and thus not cannot be comparable with $\varphi^{j+1}(x)$ because in the embedding $\varphi^{j}$, $x$ has the unique neighbor $y$.
@@
For every tree poset $P$ of height $k$, positive integers $q,\ell$ with $q\geq k$, and $\varepsilon>0$, there exists $\gamma=\gamma(\varepsilon,q,\ell)$ such that the following holds for any $\mathrm{r}:P\rightarrow[q]$ a poset homomorphism. Let $\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$ be an $\ell$-gapped family such that for all $\mathcal{S}\subseteq\mathcal{F}$ with $|\mathcal{S}|=N$ we have $\mu(\mathcal{F}-\mathcal{S})\geq q-1+\varepsilon$. Then, the number of induced copies of $P$ in $\mathcal{F}$ is at least
@@
Since $F_{1}$ must be a member of the $\mathrm{r}(x_{1})$th layer of $\mathcal{M}_{n,q}$, the number of choices for $F_{1}$ is at most $n\choose n/2$}. Given that $F_{1},\ldots,F_{j-1}$ have been selected, let $x_{i}$ be the parent of $x_{j}$. If $x_{i}<x_{j}$, then $F_{j}$ must be a set containing $F_{i}$ together with $\mathrm{r}(x_{i})-\mathrm{r}(x_{j})$ additional elements from $[n]$, and the number of such sets is at most $n^{\mathrm{r}(x_{i})-\mathrm{r}(x_{j})}$. Similarly if $x_{j}<x_{i}$ then the number of choices for $F_{j}$ is at most $n^{\mathrm{r}(x_{j})-\mathrm{r}(x_{i})}$. Multiplying the number of choices for each step gives the total result, since each edge of $H(P)$ is counted exactly once by some $n^{|\mathrm{r}(x_{i})-\mathrm{r}(x_{j})|}$ term. ∎
@@
Using Chernoff bounds, it is standard [8, 21] to show that the number of sets $F\in\mathcal{B}_{n}$ with $||F|-n/2|>2\sqrt{n\ln{n}}$ is $o({n\choose n/2})$. Thus, we may assume $\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$.
@@ 5.3 Balanced Supersaturation
For every tree poset $P$ with height $k$, real number $\varepsilon>0$, and integer $\ell$; there exists $\delta=\delta(\varepsilon,P,\ell)$ such that the following holds. Let $n$ be sufficiently large and $\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$ satisfy $|\mathcal{F}|\geq(k-1+\varepsilon)\binom{n}{n/2}$, and suppose $\mathcal{F}$ is $\ell$-gapped. Then there is a collection $\mathcal{H}$ of induced copies of $P$ from $\mathcal{F}$ satisfying
@@
    $\Delta_{j}(\mathcal{H})\leq(\delta n^{\ell})^{|P|-j}$ for all $1\leq j\leq|P|$.
@@
    For all $1\leq j\leq|P|$, $\Delta_{j}(\mathcal{H})\leq(\delta n^{\ell})^{|P|-j},$
then there exists an induced copy $P^{\prime}$ of $P$ not in $\mathcal{H}$ such that $\mathcal{H}^{\prime}=\mathcal{H}\cup\{P^{\prime}\}$ satisfies
@@ 6.1 Hypergraph Containers
[…]
@@
Let $M>0,s>0$ and $0<\delta<1$. For any sequence $(a_{1},a_{2},\dots,a_{m})$ of real numbers summing up to $s$ such that $1\leq a_{j}\leq(1-\delta)^{j}M$ for each $j\in[m]$, we have
@@ 6.2 Applications of Container Lemma
Let $k\geq 2$ and $\varepsilon>0$ be arbitrary, and let $K=K(\varepsilon,P)>0$ and $(\mathcal{T},g,C)$ be the constant and fingerprint given by Theorem 6.4 applied to $\mathcal{G}_{P}$ with $\varepsilon,P$. Let $n\in\mathbb{N}$ be sufficiently large, and note that we may assume $pn\geq K\varepsilon^{-1}$ since $pn\rightarrow\infty$. Let $\mathcal{P}(n,p)$ be the uniformly random subset of $\mathcal{B}_{n}$, where each set survives with probability $p$ . Suppose $I\subseteq\mathcal{P}(n,p)$ is an induced $P$-free subset (or equivalently an independent set of $\mathcal{G}_{P}$) of size at least $(k-1+3\varepsilon)p\binom{n}{n/2}$. Then, it follows that $\widehat{g(I)}\subseteq\mathcal{P}(n,p)$ and